Three pure compounds form when 1.50 g samples of element X combine with, respectively, 0.708 g, 0.945 g and 1.184 g of element Z. The first compound has the formula X2Z3. Find the empirical formulas of other two compounds.


1
Expert's answer
2021-10-30T01:15:54-0400

Mass of X in 1st compound = 1.0 g

Mass of Z in 1st compound = 0.472 g


The emperical formula of first compound = X2Z3

That means 1.0 g of X is equivalent to 2 moles of X

<=> 1 mole of X = 0.5 g of X.


0.472 g of Z is equivalent to 3 moles of Z

<=> 1 mole of Z = 0.15733 g of X



----


2nd compound


Mass of X = 1.0 g

So moles of X = 1.0/0.5 = 2

The mass of Z = 0.420 g

Number of moles of Z = 0.420/0.15733 = 2.6695 moles


Ratio of moles of X to Z = 2:2.6695 = 3:4


Answer: Emperical formula of second compound = X3Z4


----


3rd compound


Mass of X = 1.0 g

So moles of X = 1/0.5 = 2 moles

The mass of Z = 0.236 g

Number of moles of Z = 0.236/0.15733 = 1.5 moles


Ratio of moles of X to Z = 2:1.5 = 4:3


Emperical formula of third compound = X4Z3